OpenStudy (anonymous):

Evaluate 9x2y-2 for x = –3 and y = 2

OpenStudy (anonymous):

9(-3)*2*(2)-2 (-27)*4-2 -108-2 -110

OpenStudy (anonymous):

thats not an answer choice..

OpenStudy (anonymous):

9(-3)*2*(2)-2 = (-27)*4-2 = -108-2 = -110 sorry it is like the above one

OpenStudy (anonymous):

9(-3)*2*(2)-2

OpenStudy (anonymous):

= (-27)*4-2

OpenStudy (anonymous):

= -108-2

OpenStudy (anonymous):

= -110